Solve:
3x^2+3x-90=0
Please show me the step by steps to get the answer. Thanks

Answer by stanbon(75887)

Solve:
3x^2+3x-90=0
-----
Divide thru by 3 to get:
x^2 + x - 30 = 0
---
Factor:
(x+6)(x-5) = 0
---
x = -6 or x = 5
=======================
Cheers,
Stan H.
============